OpenStudy (anonymous):

How many points of inflection does the function F(x)=(pi/3)^(x^3 - 8) have?

OpenStudy (anonymous):

you will have to differentiate twice and set it equal to 0

OpenStudy (anonymous):

if you just need the number of inflection points, you can safely ignore the \[\frac{\pi}{3}\] out front. \[y=x^3-8\] looks just like \[y=x^3\]translated down 8 units. since \[y=x^3\] has one inflection point at (0,0) \[y=x^3-8\] also has one inflection point, this one at (0,-8)

OpenStudy (asnaseer):

unless your function is:\[F(x)=(\frac{\pi}{3})^{x^3-8}\]

OpenStudy (asnaseer):

is it the one I displayed above or is it:\[F(x)=\frac{\pi}{3}(x^3-8)\]

OpenStudy (anonymous):

No the function is F(x)=\[(\pi/3)^{x^{3}-8}\]

OpenStudy (asnaseer):

then I you cannot follow what @satellite73 showed as I think he assumed the other form of F(x). As @imranmeah91 said, what you need to do is: find the values for x that make F''(x) = 0

OpenStudy (asnaseer):

you can use wolfram to help you: http://www2.wolframalpha.com/input/?i=inflection+points+%28pi%2F3%29%5E%28x%5E3-8%29

OpenStudy (anonymous):

Thank You =)

OpenStudy (anonymous):

Nice job asnaseer

OpenStudy (anonymous):

whoa did i ever read that wrong. sorry!

OpenStudy (anonymous):

all wrong again

OpenStudy (asnaseer):

no worries @satellite73 - after all - to err is to be human...